OpenStudy (anonymous):

How do you solve Bernoulli differential equation y'-5y=3(e^t)y^2 with the conditions of y(0)=5? Please help.

OpenStudy (abb0t):

Use the substitution \(w=y^{1-n}\) to convert it into a differential euqation in terms of "w"

OpenStudy (abb0t):

However, you need to get it into standard form first! So divide everything by \(y^2\)! Essentially, you're going to massage the differential equation into a linear first oder differential equation.

OpenStudy (anonymous):

So I get y^-2(y')-5y^-1=3e^5t

OpenStudy (anonymous):

And then I have to substitute w? I'm still a bit confused...

OpenStudy (abb0t):

You have: \(\frac{1}{y^2} y'-\frac{5}{y}=e^t\) now use substitution, \(w=y^{-1}\) and find \(w'\) you should now see that you get a nice linear differential.

OpenStudy (anonymous):

Yeah, so I get -w'-5w=e^t. Wait, what happened to the 3 and the 5th power e?

OpenStudy (abb0t):

Oops, forgot about it, but it's still there. Lol. It's just a constant.

OpenStudy (anonymous):

Oh ok. So once I get w an the derivative of it and substitute it into the function. Then do I find the integrating factor?

OpenStudy (abb0t):

Yep! yOU SHOuld know what to do from here. I think you learned how to solve linear differential eequation from day 1.

OpenStudy (anonymous):

umm just a quick question. The integrating factor for this turns out to be e^(-5t) right?

OpenStudy (abb0t):

\(u(x) = e^{\int 5dt}\)

OpenStudy (anonymous):

I ended up getting 1/((3/10)-(1/10)e^(-10t)). Does that look correct?

OpenStudy (abb0t):

i can't even read that. Lol

OpenStudy (anonymous):

It looks like this.

OpenStudy (abb0t):

Is that your final solution?

OpenStudy (abb0t):

That doesn't look right to me because you're missing the constant in your solution...

OpenStudy (anonymous):

Oh I solved for the constant by using the given y(0)=5

OpenStudy (anonymous):

And the answer I gave is what it looks like after having solved for C.

OpenStudy (abb0t):

No, still wrong. Lol. If your initial condition is y(0)=5....also your exponential term is wrong. How did you get -10?

OpenStudy (anonymous):

I multiplied the integrating factor of e^(5t) to the equation with the substituted w. I think I may have done it incorrectly.

OpenStudy (abb0t):

\( \large\int [e^{5t}w]' = \int -3e^{6t}dt\)

OpenStudy (anonymous):

I got this.

OpenStudy (abb0t):

My math could be off or could be yours. But either way, close enough I guess. Lol.